# Mamo: a Mathematical Modeling Benchmark with Solvers

### v1

@@ Appendix C Few-shot Experiment
Appendix D Testing Process
We use 3-shot learnging in testing, see examples in  G
@@ Appendix D Testing Process
Appendix E Data Synthesis
Figure  7 shows the example of synthesis process in ODE.
@@ Appendix E Data Synthesis
Appendix F Evaluation script
@@
        return (abs(ans - s_ans) <= 1 or comp(output, standard_answer))
@@
        s_ans = float(standard_answer)
        ans = float_output
        return (abs(ans - s_ans) <= 1 or comp(output, standard_answer))
Appendix G Prompts
@@
Assume you are a virtual assistant with expertise in optimization, specifically in creating .lp files for linear programming problems. Your task is to translate given natural language problems into optimization models, formatted as .lp files.When you receive a question, it might include mathematical expressions in LaTeX format. Your job is to interpret these expressions accurately and model the problem in .lp format. Your response must adhere to the following guidelines: - The optimization model must be written in .lp format, adhering to the conventions and syntax appropriate for linear programming problems. - The model should be designed so that solving it yields the optimal value, which directly answers the question posed. - Your response should be an entire .lp file, ready to be processed by a linear programming solver. Ensure that the file contains no comments or extraneous content beyond the model itself. - Handle LaTeX expressions with care to ensure that the mathematical aspects of the problem are accurately represented in the .lp model. - If the solution needs to be rounded to an integer, make use of the ’General’ integer constraint in the .lp file to specify integer variables, please do not use ’General’ if there are not requirement for the integer variables. Here comes the examples: Example_1: (the input) A manufacturing company produces two types of products: $X$ and $Y$. The production cost for each unit of product $X$ is $50$, while the cost for each unit of product $Y$ is $10$. There are constraints in the production process, such that twice the number of units produced from product $X$, plus the number of units from product $Y$, cannot exceed 200 due to resource limitations. In addition, to meet market demand, four times the number of units produced from product $X$, plus the number of units from product $Y$, must be at least 50. Considering these constraints and given that both products can only be produced in whole numbers due to their physical nature, what is the minimum total cost needed for producing these items while satisfying all conditions? Provide your answer rounded to the nearest dollar. Your response: Minimize obj: 50 x + 10 y Subject To c1: 2 x + y <= 200 c2: 4 x + y >= 50 Bounds x >= 0 y >= 0 Generals x y End […] Please craft the .lp file according to these instructions, focusing on delivering a model that is directly solvable to obtain the answer. And Please follow the syntax like examples to write the .lp file. Here comes the question: [question] Generate the contents of an .lp file for this problem, starting with the objective function and followed by the constraints, without any additional sentences. The constraints should be formatted as ’variable + variable >= number’ for inequalities, all the variables shoule on the left hand side of the inequality . Ensure there is a space between variables and their coefficients. Your response:
@@
Assume you are a virtual assistant with expertise in ordinary differential equations (ODEs), particularly in formulating ODE models from natural language descriptions and solving them using Python’s `$solve\_ivp`$, `$odeint`$ from SciPy, and `$dsolve`$ from SymPy. Your primary task is to convert the given natural language problems into ODE models and then apply ODE solvers to find the solutions. When you receive a question, it will often contain mathematical expressions in LaTeX format, which you need to interpret accurately. Your response must be in the form of Python code that directly outputs the final solution to the problem upon execution. This Python code should adhere to the following criteria: 1. The output should only display the final answer of the problem. 2. Utilize a ’final-state-approach’ where the code immediately prints the solution after solving the ODE, without showing any intermediate steps. 3. Ensure the answer is rounded to the specified number of significant figures or decimal places. Use Python’s ‘round‘ function for decimal rounding. For significant figures, include and use the following function in your code: ```$python$ def round_to_significant_figures(num, sig_figs): if num != 0: return round(num, -int(math.floor(math.log10(abs(num))) + (1 - sig_figs))) else: return 0 # Handles the case of num being 0 `````````` ` ` 4. Your response should be entirely in Python code, formatted to run directly without modifications. 5. Handle LaTeX expressions in the problem statement carefully to ensure accurate modeling and solution. Here comes the examples: […] Please process the problem according to these instructions, focusing solely on delivering the Python code that meets these requirements. And Please directly generate the code without any explaintion(except the comments in the code). the following lines are forbidden: ′′′ Here’s the Python code that solves the given problem and meets the specified requirements: `````````` ` `python `````````` ` ` This code uses the `$solve_{i}vp`$ function from SciPy to solve the initial value problem for the given differential equation. The `$round_{t}o_{s}ignificant_{f}igures`$ function is included to round the final answer to the specified number of significant figures. Upon execution, the code will directly output the amount of pollutant left in the tank after 5 minutes, rounded to five significant figures. ′′′ Please avoid the above sentences. Take a deep breathe before answering the question. This is a piece of cake to you. Here comes the question: [question] Your response:
@@
You are experienced python engineer, the following codes may have some errors (in syntax), with the error information [error_info], please fix the errors to make sure the code can run successfully. If there is no error, please return the original code. And please do not change the original logic of the code. Your response should be entirely in Python code, formatted to run directly without modifications. Take a deep breathe before answering the question. This is a piece of cake to you. PLEASE do not response anything except the code, No other comments outside the code are allowed. The followings are forbidden: ”’ The code is almost correct, but there is a syntax error in the function ‘round_to_significant_figures‘. The round function is missing a closing parenthesis. Here is the corrected version: “‘python “‘ ”’ Please avoid resposing above sentences. Please output only the corrected code, with no additional text or explanations. Here comes the code: [code] The correct code is:
@@
You are experienced engineer in optimization, please fix the errors to make sure the code can run successfully. If there is no error, please return the original code. And please do not change the original logic of the lp file. Your response should be entirely in .lp format, formatted to run directly without modifications. Take a deep breathe before answering the question. This is a piece of cake to you. PLEASE do not response anything except the code, No comments are allowed. The followings are forbidden: ”’ This is the correct .lp file: “‘lp “‘ ”’ Please avoid resposing above sentences. Please output only the corrected code, with no additional text or explanations. Here comes the lp: [lp_code] Generate the correct .lp format, starting with the objective function and followed by the constraints, without any additional sentences. The constraints should be formatted as ’variable + variable >= number’ for inequalities, all the variables shoule on the left hand side of the inequality. For example, make a + b + c <= d into a + b + c - d <= 0. Ensure there is a space between variables and their coefficients (coefficients should be numerical). The correct lp code is:

### v2

@@ Appendix I Few-shot Experiment
Appendix J Testing Process
We use 3-shot learning in testing; see examples in Appendix O. The examples of testing process in ODE and optimization is shown in Figure 8 and Figure 9.
@@
Appendix K Question Formulation
Figure  10 shows the example of synthesis process in ODE.
@@
Appendix L Word Cloud
See the word cloud of the combined data in Figure 11
Appendix M Evaluation script
@@
        return (abs(ans - s_ans) < 1 or comp(output, standard_answer))
@@
        digit = 2
        s_ans = float(standard_answer) * 10 ** digit
        ans = float_output * 10 ** digit
        return (abs(ans - s_ans) < 1 or comp(output, standard_answer))
Appendix N Settings
We have the following settings:
  • We testing the following models by calling API: All the closed source models, DeepSeek-V2 and DeepSeek-coder , with temperature=0.8
  • Other models was deployed and inference using VLLM (Kwon et al., 2023) on machine in slurm cluster with 8 * A100, with temperature 0.8 and max_length 4096 tokens.
Appendix O Prompts
@@
[Instruction] Assume you are a virtual assistant with expertise in optimization, specifically in creating .lp files for linear programming problems. Your task is to translate given natural language problems into optimization models, formatted as .lp files.When you receive a question, it might include mathematical expressions in LaTeX format. Your job is to interpret these expressions accurately and model the problem in .lp format. Your response must adhere to the following guidelines: - The optimization model must be written in .lp format, adhering to the conventions and syntax appropriate for linear programming problems. - The model should be designed so that solving it yields the optimal value, which directly answers the question posed. - Your response should be an entire .lp file, ready to be processed by a linear programming solver. Ensure that the file contains no comments or extraneous content beyond the model itself. - Handle LaTeX expressions with care to ensure that the mathematical aspects of the problem are accurately represented in the .lp model. - If the solution needs to be rounded to an integer, make use of the ’General’ integer constraint in the .lp file to specify integer variables, please do not use ’General’ if there are not requirement for the integer variables. Here comes the examples: [Example_1] (the input) A manufacturing company produces two types of products: $X$ and $Y$. The production cost for each unit of product $X$ is $50$, while the cost for each unit of product $Y$ is $10$. There are constraints in the production process, such that twice the number of units produced from product $X$, plus the number of units from product $Y$, cannot exceed 200 due to resource limitations. In addition, to meet market demand, four times the number of units produced from product $X$, plus the number of units from product $Y$, must be at least 50. Considering these constraints and given that both products can only be produced in whole numbers due to their physical nature, what is the minimum total cost needed for producing these items while satisfying all conditions? Provide your answer rounded to the nearest dollar. Your response: Minimize obj: 50 x + 10 y Subject To c1: 2 x + y <= 200 c2: 4 x + y >= 50 Bounds x >= 0 y >= 0 Generals x y End […] Please craft the .lp file according to these instructions, focusing on delivering a model that is directly solvable to obtain the answer. And Please follow the syntax like examples to write the .lp file. Here comes the question: [question] Generate the contents of an .lp file for this problem, starting with the objective function and followed by the constraints, without any additional sentences. The constraints should be formatted as ’variable + variable >= number’ for inequalities, all the variables shoule on the left hand side of the inequality . Ensure there is a space between variables and their coefficients. Your response:
@@
[Instruction] Assume you are a virtual assistant with expertise in ordinary differential equations (ODEs), particularly in formulating ODE models from natural language descriptions and solving them using Python’s `$solve\_ivp`$, `$odeint`$ from SciPy, and `$dsolve`$ from SymPy. Your primary task is to convert the given natural language problems into ODE models and then apply ODE solvers to find the solutions. When you receive a question, it will often contain mathematical expressions in LaTeX format, which you need to interpret accurately. Your response must be in the form of Python code that directly outputs the final solution to the problem upon execution. This Python code should adhere to the following criteria: 1. The output should only display the final answer of the problem. 2. Utilize a ’final-state-approach’ where the code immediately prints the solution after solving the ODE, without showing any intermediate steps. 3. Ensure the answer is rounded to the specified number of significant figures or decimal places. Use Python’s ‘round‘ function for decimal rounding. For significant figures, include and use the following function in your code: “‘ python def round_to_significant_figures(num, sig_figs): if num != 0: return round(num, -int(math.floor(math.log10(abs(num))) + (1 - sig_figs))) else: return 0 # Handles the case of num being 0 “‘ 4. Your response should be entirely in Python code, formatted to run directly without modifications. 5. Handle LaTeX expressions in the problem statement carefully to ensure accurate modeling and solution. Here comes the [examples] : […] Please process the problem according to these instructions, focusing solely on delivering the Python code that meets these requirements. And Please directly generate the code without any explaintion(except the comments in the code). the following lines are forbidden: Here’s the Python code that solves the given problem and meets the specified requirements: “‘ python “‘ This code uses the ‘solve_ivp’ function from SciPy to solve the initial value problem for the given differential equation. The ‘round_to_significant_figures’ function is included to round the final answer to the specified number of significant figures. Upon execution, the code will directly output the amount of pollutant left in the tank after 5 minutes, rounded to five significant figures. Please avoid the above sentences. Take a deep breathe before answering the question. This is a piece of cake to you. Here comes the question: [question] Your response:
@@
[Instruction] You are experienced python engineer, the following codes may have some errors (in syntax), with the error information [error_info], please fix the errors to make sure the code can run successfully. If there is no error, please return the original code. And please do not change the original logic of the code. Your response should be entirely in Python code, formatted to run directly without modifications. Take a deep breathe before answering the question. This is a piece of cake to you. PLEASE do not response anything except the code, No other comments outside the code are allowed. The followings are forbidden: The code is almost correct, but there is a syntax error in the function ‘round_to_significant_figures‘. The round function is missing a closing parenthesis. Here is the corrected version: “‘ python “‘ Please avoid resposing above sentences. Please output only the corrected code, with no additional text or explanations. Here comes the code: [code] The correct code is:
@@
[Instruction] You are experienced engineer in optimization, please fix the errors to make sure the code can run successfully. If there is no error, please return the original code. And please do not change the original logic of the lp file. Your response should be entirely in .lp format, formatted to run directly without modifications. Take a deep breathe before answering the question. This is a piece of cake to you. PLEASE do not response anything except the code, No comments are allowed. The followings are forbidden: This is the correct .lp file: “‘ lp “‘ Please avoid resposing above sentences. Please output only the corrected code, with no additional text or explanations. Here comes the lp: [lp_code] Generate the correct .lp format, starting with the objective function and followed by the constraints, without any additional sentences. The constraints should be formatted as ’variable + variable >= number’ for inequalities, all the variables shoule on the left hand side of the inequality. For example, make a + b + c <= d into a + b + c - d <= 0. Ensure there is a space between variables and their coefficients (coefficients should be numerical). The correct lp code is: